OpenStudy (anonymous):

Find the limit of the function algebraically. limx->-6 x2 -36/x + 6

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

factor cancel replace \(x\) by \(-6\)

OpenStudy (anonymous):

i can walk you through it if you like , but those are the steps

OpenStudy (anonymous):

Okay how do I factor? I can do the rest. I think.

OpenStudy (anonymous):

factor \[x^2-36=(x+6)(x-6)\]

OpenStudy (anonymous):

\(\Large\frac{x^2-36}{x+6}=\frac{(x+6)(x-6)}{(x+6)}\)

OpenStudy (anonymous):

i.e. the difference of two squares

OpenStudy (anonymous):

then cancel, get \(x-6\) then replace \(x\) by \(-6\)

OpenStudy (anonymous):

I think there's a -6 though as the coefficient of x^2. At least that's how I am interpreting what she originally wrote.

OpenStudy (anonymous):

No, It is just asking for the one sided limit coming for the left at x=-6

OpenStudy (anonymous):

Ahh I see okay, thanks for the clarification. My mistake.

OpenStudy (anonymous):

here is all the steps, written in one line \[\lim_{x\to -6}\frac{x^2-36}{x+6}=\lim_{x\to -6}\frac{(x+6)(x-6)}{x+6}\\ =\lim_{x\to -6}x-6=-6-6=-12\]

OpenStudy (anonymous):

Aahhhh thank you for the clarification

OpenStudy (anonymous):

it is not a "one sided limit" it is just the limit hope the steps are clear, very many work exactly this way

OpenStudy (anonymous):

Oh yeah sorry I also read it wrong because of the -> nice work @satellite73

OpenStudy (anonymous):

you want to try another or are you done?

OpenStudy (anonymous):

Ummm can you help with this one? Find the center, vertices, and foci of the ellipse with equation 5x^2 + 8y^2 = 40.

OpenStudy (anonymous):

sure but this is totally different

OpenStudy (anonymous):

I know. Sorry.

OpenStudy (anonymous):

it is not that hard though start by dividing by 40 to get \[\frac{x^2}{8}+\frac{y^2}{5}=1\]

OpenStudy (anonymous):

Okay what's next?

OpenStudy (anonymous):

then you are really almost done is it clear that the centre is \((0,0)\) ?

OpenStudy (anonymous):

Yes I see that

OpenStudy (anonymous):

ok and it is also clear that the larger number is under the \(x^2\) term, so it looks like this 

OpenStudy (anonymous):

Okay I'm with you so far

OpenStudy (anonymous):

general form is \[\frac{x^2}{a^2}+\frac{y^2}{b^2}=1\] if the center is \((0,0)\) that makes \(a^2=8\) and so \(a=\sqrt8=2\sqrt2\)

OpenStudy (anonymous):

you need to know what it looks like to see that the foci are \(a\) units to the RIGHT and LEFT of the center so the foci are \((-2\sqrt2,0)\) and \((2\sqrt2,0)\)

OpenStudy (anonymous):

whoa scratch that!!!

OpenStudy (anonymous):

all of what i said there was true, but it is for the VERTICES not the foci my mistake

OpenStudy (anonymous):

vertices are \((-2\sqrt2,0)\) and \((2\sqrt2,0)\)

OpenStudy (anonymous):

damn i am full of typos let me know if you are confused if not we can go on and find the foci it is pretty easy really

OpenStudy (anonymous):

Okay so the foci are \[(-\sqrt{3} ,0) (\sqrt{3}, 0)\]

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

Can we do another?

OpenStudy (anonymous):

lol sure

OpenStudy (anonymous):

hope i can go easier on the typos

OpenStudy (anonymous):

Find the vertices and foci of the hyperbola with equation (x + 2)^2/144 - (y - 4)^2/81= 1

OpenStudy (anonymous):

Sorry, I realized I never hit post

OpenStudy (anonymous):

k at least this one is in the right form to start you got the center?

OpenStudy (anonymous):

\[\frac{(x+2)^2}{12^2}-\frac{(y-1)^2}{9^2}=1\]

OpenStudy (anonymous):

Okay so the center is (0,0)

OpenStudy (anonymous):

right?

OpenStudy (anonymous):

oh no

OpenStudy (anonymous):

oh...I'm confused

OpenStudy (anonymous):

\[\frac{(x+2)^2}{12^2}-\frac{(y-1)^2}{9^2}=1\] \[\frac{(x-h)^2}{a^2}-\frac{(y-k)^2}{b^2}=1\] center is \((h,k)\)

OpenStudy (anonymous):

they have shifted it is not \(x^2\) and \(y^2\) up top your center for \[\frac{(x+2)^2}{12^2}-\frac{(y-1)^2}{9^2}=1\] is \((-2,1)\)

OpenStudy (anonymous):

ohhhhhh okay

OpenStudy (anonymous):

ok we need that first we also need to know what it looks like